### v2

# preprint: IFT-UAM/CSIC-25-123

Abstract
We establish a direct correspondence between Krylov and Nielsen complexity by choosing the Krylov basis to be part of the elementary gate set of Nielsen geometry and selecting a Nielsen complexity metric compatible with the Krylov metric. Up to normalization, the Krylov complexity of a Hermitian operator then equals the length squared of a straight-line trajectory on the manifold of unitaries that connects the identity operator with a precursor operator. The corresponding length provides an upper bound on Nielsen complexity that saturates whenever the straight line is a minimal geodesic. While for general systems we can only establish saturation in the limit of small precursors, we provide evidence that in the Sachdev–Ye–Kitaev (SYK) model there is a precise correspondence between Krylov complexity and (the square of) Nielsen complexity for a finite range of precursors.
preprint: IFT-UAM/CSIC-25-123
Introduction. Quantum complexity has become a central concept at the intersection of quantum information, condensed-matter theory, and high-energy physics. One of the most influential approaches is Nielsen complexity [Nielsen:2005mkt, Nielsen:2006cea, Dowling:2006tnk], which reformulates circuit complexity as a geometric problem. In this framework, a circuit corresponds to a curve on the unitary group manifold, and the complexity is given by the length of the minimal geodesic connecting the chosen unitary operator with the identity operator [Jefferson:2017sdb, Khan:2018rzm, Hackl:2018ptj, Chapman:2018hou, Bhattacharyya:2018bbv, Guo:2018kzl, Bernamonti:2019zyy, Bernamonti:2020bcf, Caceres:2019pgf, Ruan:2020vze, Haque:2024ldr]. A complementary measure, Krylov complexity [Parker:2018yvk], quantifies operator growth under Heisenberg evolution. Starting from a seed operator, one constructs an orthonormal Krylov basis via the Lanczos algorithm, thereby mapping the dynamics to a one-dimensional hopping problem along the Krylov chain. Krylov complexity directly measures the spread of the initial operator in Krylov space and provides a quantitative probe of operator growth, scrambling, and chaos [Barbon:2019wsy, Avdoshkin:2019trj, Rabinovici:2020ryf, Jian:2020qpp, Dymarsky:2021bjq, Hornedal:2022pkc, Balasubramanian:2022tpr, Erdmenger:2023wjg, Caputa:2024vrn, Baggioli:2024wbz, Craps:2024suj, Huh:2024ytz]. For comprehensive reviews, see [Nandy:2024htc, Baiguera:2025dkc, Rabinovici:2025otw].
Despite their different origins, Nielsen complexity and Krylov complexity both have deep ties to holography. Nielsen’s geometric framework has motivated holographic proposals linking circuit complexity to geometric quantities that capture the growth of the black hole interior [Susskind:2014rva, Stanford:2014jda, Brown:2015bva, Brown:2015lvg, Cai:2016xho, Belin:2021bga, Belin:2022xmt, Carrasco:2023fcj, Jorstad:2023kmq, Jiang:2023jti, Caceres:2023ziv, Myers:2024vve, Arean:2024pzo, Jiang:2025qai, Miyaji:2025yvm, Miyaji:2025jxy, Caceres:2025myu]. Meanwhile, in models of two-dimensional gravity, Krylov state complexity provides a concrete realization of the proposed correspondence between complexity and size of black hole interiors [Lin:2022rbf, Rabinovici:2023yex, Heller:2024ldz, Balasubramanian:2024lqk, Aguilar-Gutierrez:2024nau, Xu:2024gfm, Heller:2025ddj, Ambrosini:2025hvo, Fu:2025kkh]. Although the possible relation between Nielsen complexity and Krylov complexity has attracted significant recent interest [Lv:2023jbv, Caputa:2021sib, Craps:2023ivc, Aguilar-Gutierrez:2023nyk], a direct and general correspondence remains elusive. These developments raise a central question: are Nielsen complexity and Krylov complexity fundamentally related? The aim of this work is to provide such an explicit correspondence.
Mapping Krylov to Nielsen complexity. Krylov complexity characterizes the dynamical growth of a Hermitian operator under time evolution. We begin with a normalized initial operator $\mathcal{O}(t=0)\equiv\mathcal{O}_{0}$, defined with respect to a chosen inner product, such as the infinite-temperature Wightman inner product,
$(\mathcal{O}|\mathcal{O}^{\prime}):=\frac{1}{\operatorname{Tr}[1]}\operatorname{Tr}\left[\mathcal{O}^{\dagger}\mathcal{O}^{\prime}\right]\,.$ (1)
The Heisenberg evolution of the operator under a time-independent Hamiltonian $H$, $\mathcal{O}(t)=e^{iHt}\mathcal{O}(0)e^{-iHt}$, is naturally formulated in operator space using the Liouvillian superoperator $\mathcal{L}\equiv[H,\cdot]$. Successive applications of $\mathcal{L}$ to the initial operator generate the Krylov subspace $\mathcal{K}=\mathrm{span}\{\mathcal{O}_{0},i\mathcal{L}\mathcal{O}_{0},(i\mathcal{L})^{2}\mathcal{O}_{0},\dots\}$. An orthonormal basis $\{|\mathcal{O}_{n})\}$ for this subspace can be constructed recursively via the Lanczos algorithm, i.e., by applying the Gram–Schmidt procedure to the Krylov sequence $\mathcal{K}$. Using this orthonormal and Hermitian basis, the time-evolved operator can be expanded as
$|\mathcal{O}(t))\equiv e^{i\mathcal{L}t}|\mathcal{O}(0))=\sum_{n}\varphi_{n}(t)|\mathcal{O}_{n})\,$, (2)
with the initial condition $\varphi_{n}(0)=\delta_{n0}$  [^1] Unlike some conventions in the literature, we absorb the factor $i^{n}$ into the basis $|\mathcal{O}_{n})$ so that each element is manifestly Hermitian.. Krylov complexity is thus defined as [Parker:2018yvk, Balasubramanian:2022tpr, Chen:2024imd]
$C_{\textrm{\tiny K}}(\mathcal{O}(t)):=\sum_{m,n}f_{mn}\,\varphi_{m}(t)\varphi_{n}(t)\,$, (3)
where the positive matrix $f_{nm}$, referred to as the Krylov metric, sets the weights for operator growth. A standard choice is $f_{mn}=f(n)\,\delta_{mn}$ with $f(n)=n$  [^2] Since $f(0)=0$, the matrix is only positive semidefinite; we set $f(0)=\delta$ and take $\delta\to 0$ at the end. More generally, any monotonically increasing $f(n)$ is admissible [Balasubramanian:2022tpr]., yielding a natural measure of spreading along the Krylov chain. Another interesting choice, consistent with the conjecture that the growth rate of Nielsen complexity should be proportional to operator size [Susskind:2014jwa, Roberts:2014isa], is the Krylov metric $f_{mn}=n^{2}\delta_{mn}$, as discussed in part A of the Supplemental Material.
In contrast, Nielsen complexity is defined for unitary transformations, which mimic sequences of elementary unitary gates. The central idea of Nielsen’s geometric formalism is to interpret the construction of a unitary operator $U(s)$ as a smooth trajectory on the manifold of unitaries, with control Hamiltonian (Hermitian velocity)
$H_{\mathrm{c}}(s)=i\partial_{s}UU^{-1}\,,\,\,H_{\mathrm{c}}(s)\equiv Y^{I}(s)T_{I}\,.$ (4)
This Hamiltonian is expanded in a basis of Hermitian generators $\{T_{I}\}$ representing the elementary gates, with control functions $Y^{I}(s)$ playing the role of tangent vectors along the trajectory. Taking the initial point as the identity, $U(0)=\mathbbm{1}$, and the endpoint as the target unitary, $U(1)=U_{\textrm{\tiny T}}$, Nielsen complexity is then defined as the minimal length (also known as minimal “cost”) among all admissible paths,
$C_{\textrm{\tiny N}}(U_{\textrm{\tiny T}})=\min_{Y^{I}(s)}\int_{0}^{1}\,\sqrt{\sum_{I,J}G_{IJ}Y^{I}(s)Y^{J}(s)}\,ds\,$, (5)
where the positive, right-invariant “complexity metric” (or “cost function”) $G_{IJ}$ encodes the weights assigned to each gate. As a result, minimizing the cost is equivalent to solving the geodesic equation (or Euler–Arnold equation).
A natural bridge between these two notions of complexity can be established by focusing on a special class of unitaries known as precursors [Susskind:2013lpa], defined as
$U_{\textrm{\tiny T}}=e^{iHt}e^{-iz\mathcal{O}(0)}e^{-iHt}=e^{-iz\mathcal{O}(t)}\,$, (6)
where $\mathcal{O}(t)$ is normalized using (1) and the positive parameter $z$ is the norm of the Hermitian generator $z\,\mathcal{O}(t)$ of $U_{\textrm{\tiny T}}$. Equation (6) links the time-evolved Hermitian operator $\mathcal{O}(t)$ that enters Krylov complexity to $U_{\textrm{\tiny T}}$ in Nielsen’s formalism. As we show below, the parameter $z$ will play a central role in relating the two notions of complexity. It is now straightforward to connect the identity to the target unitary via the straight-line trajectory $U(s)=e^{-isH_{\mathrm{con}}}$, which corresponds to a constant control Hamiltonian,
$H_{\mathrm{c}}(s)\rightarrow H_{\mathrm{con}}\equiv z\mathcal{O}(t)=z\sum_{n}\varphi_{n}(t)\mathcal{O}_{n}\,.$ (7)
The next step is to choose the basis of Hermitian generators $\{T_{I}\}$ used in Nielsen’s construction to contain the Krylov basis $\{|\mathcal{O}_{n})\}$, i.e., $\{\mathcal{O}_{n}\}\subseteq\{T_{I}\}$. A simple correspondence is then realized by identifying
$T_{n}\sim\mathcal{O}_{n}\,,\quad Y^{n}\sim z\varphi_{n}\,.$ (8)
Because the path $U(s)$ is a straight line with constant controls, its cost can be easily read off once the complexity metric is specified:
$\operatorname{Cost}(\mathcal{O}(t))=z\sqrt{\sum_{m,n}G_{mn}\,\varphi_{m}\varphi_{n}}\geq C_{\textrm{\tiny N}}(U_{\textrm{\tiny T}})\,$, (9)
which provides an upper bound on the Nielsen complexity of the precursor $U_{\textrm{\tiny T}}$. Choosing $G_{IJ}$ such that $G_{mn}=f_{mn}$ for the directions corresponding to the Krylov basis, the two measures obey
$C_{\textrm{\tiny N}}(e^{-iz\mathcal{O}(t)})\leq z\,\sqrt{C_{\textrm{\tiny K}}(\mathcal{O}(t))}\,.$ (10)
In fact, for targets close to the identity ($z\ll 1$) [Ali:2019zcj, Qu:2021ius, Qu:2022zwq], the leading-order expansion yields  [^3] Higher-order corrections in this expansion have been previously discussed in [Haque:2024ldr].
$C_{\textrm{\tiny N}}(e^{-iz\mathcal{O}(t)})=z\sqrt{C_{\textrm{\tiny K}}(\mathcal{O}(t))}+O(z^{3})\,$, (11)
explicitly linking the two notions of complexity for “small” precursor operators in general systems.
The remainder of this work addresses the question of whether the saturation of (10) extends to finite $z$, i.e., beyond the infinitesimal precursor limit. In other words, is the straight-line trajectory the shortest path from the identity operator to the target unitary $U_{\textrm{\tiny T}}$? While for generic systems a trajectory generated by a constant control Hamiltonian $H_{\rm con}$ is not a geodesic, we will find that the geodesic equation is satisfied for the interesting special case of the SYK model [Sachdev_1993, KitaevTalks]. For the inequality (10) to be saturated,
$C_{\textrm{\tiny N}}(e^{-iz\mathcal{O}(t)})=z\,\sqrt{C_{\textrm{\tiny K}}(\mathcal{O}(t))}\,$, (12)
it should moreover be checked that the geodesic is minimal.
To make the connection between the formalisms of Krylov and Nielsen complexities more explicit, we identify the inner product in both frameworks with Eq. (1) and choose the gate basis $\{T_{I}\}$ so that $(T_{I},T_{J})=\delta_{IJ}$. The geodesic equation can then be recast in the compact form [Dowling:2006tnk]
$\dot{H}_{\rm c}+i\,\mathbf{G}^{-1}\,([H_{\mathrm{c}},\mathbf{G}(H_{\mathrm{c}})])=0\,$, (13)
where $\mathbf{G}(H_{\rm c})=\sum_{I,J}G_{IJ}Y^{J}T_{I}$  [^4] The chosen trace pairing $(T_{I},T_{J})=\delta_{IJ}$ naturally identifies the dual algebra $\mathfrak{g}^{\ast}$ with the Lie algebra $\mathfrak{g}$.. It follows immediately that a constant control Hamiltonian, $\dot{H}_{\rm c}=0$, constitutes a geodesic whenever $\mathbf{G}^{-1}\!\left([H_{\mathrm{con}},\mathbf{G}(H_{\mathrm{con}})]\right)=0$.
Simplifications in the SYK model. The $q$-body SYK model consists of $N$ Majorana fermions with all-to-all random interactions [Maldacena:2016hyu]. Its Hamiltonian is given by
$H=i^{q/2}\sum_{1\leq i_{1}<\cdots<i_{q}\leq N}J_{i_{1}\cdots i_{q}}\,\psi_{i_{1}}\cdots\psi_{i_{q}}\,$, (14)
where the couplings $J_{i_{1}\cdots i_{q}}$ are drawn from a Gaussian ensemble with zero mean and variance
$\langle J_{i_{1}i_{2}\ldots i_{q}}^{2}\rangle=\frac{2^{q-1}\mathcal{J}^{2}(q-1)!}{q\,N^{q-1}}\,.$ (15)
We note that the space of linear operators on the $2^{N/2}$-dimensional Hilbert space (with an operator basis ${X_{\alpha}}\equiv{\psi_{i_{1}}\cdots\psi_{i_{k}}}$, $\alpha=1,\ldots,2^{N}$) naturally decomposes into two orthogonal sectors. Introducing the chirality operator:
$\psi^{5}=(2i)^{N/2}\psi_{1}\psi_{2}\cdots\psi_{N}\,$, (16)
and noting that $(\psi^{5})^{2}=\boldsymbol{1}$ and $\{\psi^{5},\psi_{i}\}=0$, one finds that traces of Majorana strings vanish for odd $k$, i.e., $\operatorname{Tr}(\psi_{i_{1}}\cdots\psi_{i_{k}})=0$. This property allows us to decompose the operator space into an odd sector $X_{\rm o}$ and an even sector $X_{\rm e}$, with algebraic structure given by
$X_{\rm o},X_{\rm o}]\sim X_{\rm e}\,,\,\,[X_{\rm o},X_{\rm e}]\sim X_{\rm o}\,,\,\,[X_{\rm e},X_{\rm e}]\sim X_{\rm e}\,.$ (17)
This simple fact greatly simplifies our analysis of the connection between the two complexity measures. Starting from an odd initial operator, e.g., $\mathcal{O}_{0}=\sqrt{2}\,\psi_{1}$, it is easy to see that the Krylov subspace $\mathcal{K}$ lies entirely in the odd sector ($\mathcal{O}_{n}\in X_{\rm o}$), since the SYK Hamiltonian belongs to $X_{\rm e}$. This motivates a block-diagonal ansatz for the complexity metric, which we take to be:
$\mathbf{G}=\begin{pmatrix}\mathbf{G}_{\rm{odd}}&0\\ 0&\mathbf{G}_{\rm{even}}\end{pmatrix}\,$. (18)
We now examine whether the straight-line trajectory $H_{\rm c}=H_{\rm con}\in X_{\rm o}$ satisfies the geodesic equation (13). We observe that a simple solution arises if the penalty factor is homogeneous across Krylov sub-directions: $\mathbf{G}(H_{\mathrm{con}})=\mathbf{G}_{\rm odd}(H_{\mathrm{con}})\propto H_{\mathrm{con}}\,.$ For more general penalty factors, however, the commutator $H_{\mathrm{con}},\mathbf{G}(H_{\mathrm{con}})]$ does not vanish. Nevertheless, owing to the algebraic fact that $H_{\mathrm{con}},\mathbf{G}(H_{\mathrm{con}})]\in X_{\rm e}$, substituting the constant $H_{\rm con}$ into the geodesic equation yields:
$\mathbf{G}^{-1}\,([H_{\mathrm{c}},\mathbf{G}(H_{\mathrm{c}})])\sim(G_{\rm{even}})^{-1}\,G_{\rm{odd}}\,.$ (19)
Here, the subspace indices and the structure constants of the unitary group are suppressed for simplicity. One then finds that the constant solution $H_{\rm c}=H_{\rm con}$ asymptotically satisfies the geodesic equation in the limit $(G_{\rm{even}})^{-1}\,G_{\rm{odd}}\to 0$ (equivalently, $G_{\rm even}\to\infty$ for fixed $G_{\rm odd}$). This indicates that the straight-line trajectory within the Krylov subspace becomes a geodesic in the full operator space whenever the even sector is heavily penalized, i.e., when its contribution is effectively suppressed by a large penalty factor.
Minimal geodesics in SYK. Nielsen complexity is defined as the minimal cost over all admissible paths, so it is essential to verify that the corresponding geodesic is indeed minimal, both locally and globally. A geodesic is locally length-minimizing if and only if it contains no conjugate points along its trajectory [frankel2004geometry], as illustrated in Fig. 1.
To investigate the presence of such conjugate points in the SYK example, we parameterize small perturbations around the reference trajectory as
$\delta H_{\rm c}(s)=\delta H_{\rm o}(s)+\delta H_{\rm e}(s)\,.$ (20)
Solving the perturbed geodesic equation with arbitrary penalty factors in the SYK model is generally difficult. To gain intuition, we solve it for a special case, namely homogeneous penalties in the odd and even sectors of the SYK model. Furthermore, in part B of the Supplemental Material, we solve the perturbed geodesic equation for the Berger sphere with arbitrary penalties. In all these cases, the location of the first conjugate point is bounded from below by $\pi/2$, which can be taken to suggest that also in more general situations the first conjugate point will stay a finite distance away from $z=0$.
Figure 1: Conjugate points (red dot) along the geodesic (black straight-line) are defined as the points where the perturbation satisfying the geodesic deviation equation vanishes.
So, we assume homogeneous penalties in the odd and even sectors,
$\mathbf{G}(X_{\rm o})=G_{\rm odd}\,X_{\rm o}\,,\quad\mathbf{G}(X_{\rm e})=G_{\rm even}\,X_{\rm e}\,.$ (21)
Under this simplification, the straight-line trajectory with a constant control Hamiltonian $H_{\rm con}=z\,\mathcal{O}(t)$ can be shown to satisfy the geodesic equation (13), thereby representing a geodesic in the space of unitaries. The associated geodesic deviation (Jacobi) equation governing small perturbations $\delta H_{\rm c}$ then takes the form [Balasubramanian:2019wgd]
$i\frac{d}{ds}\delta H_{\rm{o}}=(\alpha-1)[H_{\rm{con}},\delta H_{\rm{e}}],\qquad i\frac{d}{ds}\delta H_{\rm{e}}=0\,$, (22)
where we have introduced the ratio $\alpha\equiv G_{\rm even}/G_{\rm odd}$. After fixing the initial conditions, the corresponding solution of the geodesic deviation equation (i.e., the Jacobi fields) can be expressed as
$\delta H_{\rm c}(s)=\delta H_{\rm c}(0)-i(\alpha-1)\,s\,[H_{\rm{con}},\delta H_{\rm{e}}(0)]\,.$ (23)
Solutions for conjugate points in a closely related setup were analyzed in detail in Ref. [Balasubramanian:2019wgd]. The essential observation is that the first-order perturbation at the endpoint of the straight-line trajectory is given by
$(U^{-1}\delta U)|_{s=1}=\int_{0}^{1}d\tilde{s}\,e^{iH_{\rm con}\tilde{s}}\delta H_{\rm c}(\tilde{s})e^{-iH_{\rm con}\tilde{s}}\equiv\mathbf{\Delta}[\delta H_{\rm c}]\,$, (24)
which can be interpreted as the action of a superoperator $\boldsymbol{\Delta}$ on the perturbation $\delta H_{\rm c}$. The conjugate points are therefore determined by the zeros of this operator, i.e., by the condition $(U^{-1}\delta U)|_{s=1}=0$. To evaluate the integral in Eq. (24) explicitly, we expand in the eigenbasis of the Hermitian operator $\mathcal{O}(t)$, defined by $\mathcal{O}(t)\,|n\rangle=\omega_{n}|n\rangle$  [^5] The time dependence is encoded in the eigenstates as $|n\rangle\equiv|n(t)\rangle=e^{iHt}|n_{0}\rangle$, where $\mathcal{O}(0)\,|n_{0}\rangle=\omega_{n}|n_{0}\rangle$.. Assuming the initial perturbations take the form
$\delta H_{\rm c}(0)=\sum_{m,n}S_{mn}|m\rangle\langle n|\,,\quad\delta H_{\rm{e}}(0)=\sum_{m,n}M_{mn}|m\rangle\langle n|\,$,
the integration in Eq. (24) yields
[…]
where $\omega_{mn}=|\omega_{m}-\omega_{n}|$. The solution for the geodesic perturbation $\delta H_{\rm c}(s)$ in Eq. (23) is determined by an arbitrary initial perturbation spanning the operator space, i.e., $\delta H_{\rm c}(0)=\sum_{\alpha}c_{\alpha}X_{\alpha}$. The existence of a conjugate point, $\boldsymbol{\Delta}[\delta H_{\rm c}]=0$, corresponds to the vanishing of the determinant:
$0=\det\mathbf{\Delta}_{\alpha\beta}\equiv\det\left(\text{Tr}\left(X_{\alpha}\mathbf{\Delta}[X_{\beta}]\right)\right)\,.$ (25)
Figure 2: Absolute value of the determinant in Eq. (25) as a function of the precursor size $z$ for various penalty ratios $\alpha\equiv G_{\rm even}/G_{\rm odd}$. The system size of the SYK model is fixed at $N=8$.
Numerical evaluation of this determinant for the SYK model is shown in Fig. 2. Although obtaining a closed-form expression for the exact locations of the conjugate points is challenging, it is straightforward to observe that, in the special case $\alpha=1$, corresponding to the bi-invariant complexity metric, the determinant vanishes at $z=\tfrac{2\pi k}{\omega_{mn}}$ for integer $k$. The first conjugate point therefore occurs at $z_{\ast}=\tfrac{2\pi}{\omega_{\max}}$, where $\omega_{\max}$ denotes the maximum spectral gap. Choosing $\mathcal{O}(t)=\sqrt{2}\,\psi_{1}(t)$ gives $\omega_{\max}=2$, and thus $z_{\ast}=\pi$. Intuitively, reducing the penalty assigned to even-sector operators (i.e., decreasing $\alpha$) brings the conjugate points closer, whereas increasing $\alpha$ pushes them further away. This trend is confirmed by our numerical results. More importantly, the location of the first conjugate point is bounded below:
[…]
with the lower bound saturated at $\alpha=0$. In part B of the Supplemental Material, we analytically verify that this bound also holds for the Berger sphere with arbitrary penalties. We thus find that, in all models that we investigated, the straight-line trajectory (7) remains a locally minimal geodesic for any precursor operator with $z\leq\tfrac{\pi}{2}$. This motivates us to expect that, also for the SYK model with penalty schedules of interest, the straight-line trajectory (7) will be a locally minimal geodesic for a finite range of $z$.
To complete our construction, we now discuss the global minimality of geodesics. Since the unitary group $\mathrm{SU}(N)$ is compact, geodesic loops can occur: beyond a certain distance, two distinct geodesics may connect the same endpoints with equal length. Starting from the Hermitian operator $\mathcal{O}(0)=\sqrt{2}\,\psi_{1}$, whose eigenvalues are fixed at $\pm 1$, it is straightforward to see that the geodesic forms a closed loop at $z=2\pi$, since
$U_{\textrm{\tiny T}}|_{z=2\pi}=\mathbbm{1}=U(s=0)\,.$ (27)
Combining this with the local minimality condition, we conclude that the locally minimizing geodesic remains globally minimizing up to the midpoint $z=\pi$. In this regime, we establish an explicit correspondence between the Krylov complexity $C_{\textrm{\tiny K}}(\mathcal{O}(t))$ and the Nielsen complexity $C_{\textrm{\tiny N}}(e^{-iz\,\mathcal{O}(t)})$ by showing that the straight-line trajectory $U(s)=e^{-isz\,\mathcal{O}}$ is the globally minimizing geodesic underlying both measures of complexity.
Summary. In this work, we have established a direct correspondence between Krylov complexity and Nielsen complexity by embedding the Krylov basis and weight metric into the elementary gate set and complexity metric of Nielsen’s geometric approach. We showed that Krylov complexity is mapped to the square of the length of a straight-line trajectory on the manifold of unitaries, which generally serves as an upper bound on Nielsen complexity. The bound generally saturates in the limit of “small” precursors ($z\to 0$). We presented evidence that it also saturates for a finite range of $z$ in the physically interesting setting of the SYK model. Our results thus establish a concrete geometric bridge between two distinct yet complementary measures of quantum complexity.
Acknowledgements.
Acknowledgments. We are grateful to José Barbón, Hong-Yue Jiang, and Yu-Xiao Liu for valuable discussions and comments on the manuscript. Collaboration with Marine De Clerck, Oleg Evnin and Philip Hacker on related topics is also gratefully acknowledged. Work at VUB is supported by FWO-Vlaanderen project G012222N and by the VUB Research Council through the Strategic Research Program High-Energy Physics. GP is further supported by a PhD fellowship from FWO. JFP is supported by the Spanish MINECO ‘Centro de Excelencia Severo Ochoa’ program under grant SEV-2012-0249, the Comunidad de Madrid ‘Atracción de Talento’ program (ATCAM) grant 2020-T1/TIC-20495, the Spanish Research Agency via grants CEX2020-001007-S and PID2021-123017NB-I00, funded by MCIN/AEI/10.13039/501100011033, and ERDF ‘A way of making Europe.’ LCQ is funded by a scholarship from the China Scholarship Council (CSC) and acknowledges support from a short-term scientific mission grant under COST Action CA22113 THEORY-CHALLENGES. LCQ also thanks the Vrije Universiteit Brussel (VUB) for hospitality while part of this work was carried out. SMR is supported by Peking University under startup Grant No. 7101303985, and gratefully acknowledges the hospitality of the Instituto de Física Teórica (IFT UAM/CSIC) and Beijing Normal University while part of this work was carried out.
Supplemental Material
Appendix A A. SYK model and operator size
In the main text, we demonstrated that the straight-line control $H_{\rm con}=z\mathcal{O}(t)$ satisfies the geodesic equation (13) in the SYK model. Of particular interest is the large-$N$ regime with $N\gg q^{2}\gg 1$. In this limit, the $n$-th Krylov basis element $\mathcal{O}_{n}$ can be expressed as a superposition of Majorana strings of fixed length [Roberts:2018mnp, Parker:2018yvk, Bhattacharjee:2022lzy],
$\mathcal{O}_{n}=i^{n}\sum_{i_{1}<\cdots<i_{s}}c_{i_{1}\cdots i_{s}}\,\psi_{i_{1}}\cdots\psi_{i_{s}}+O\left(\frac{1}{q}\right)\,$, (S1)
where the operator size is s=n(q2)+1s=n(q-2)+1, with $n$ labeling the Krylov generation [Roberts:2018mnp, Parker:2018yvk]. Correspondingly, the associated Krylov wavefunctions admit compact analytic forms [Roberts:2018mnp, Parker:2018yvk]:
$\varphi_{n}(t)=\begin{cases}1+\dfrac{2}{q}\,\ln\text{sech}(\mathcal{J}t)+O\left(\tfrac{1}{q^{2}}\right),&n=0,\\[6.0pt] \sqrt{\dfrac{2}{nq}}\,\tanh^{n}(\mathcal{J}t)+O\left(\tfrac{1}{q^{2}}\right),&n\geq 1\,.\end{cases}$ (S2)
The average operator size then follows as
$s(t)=\sum_{n}\left[1+(q-2)n\right]\varphi_{n}^{2}(t)=\cosh(2\mathcal{J}t)+O\!\left(\tfrac{1}{q^{2}}\right)\,.$ (S3)
A particularly illuminating choice of the Krylov metric arises when adopting a quadratic form,
$f_{nm}=n^{2}\,\delta_{nm}\,$, (S4)
in contrast to the more common linear choice $f_{nm}=n\,\delta_{nm}$. This quadratic weighting leads to a Nielsen complexity of the form
$C_{\textrm{\tiny N}}(e^{-iz\mathcal{O}(t)})=z\sqrt{\sum_{n}n^{2}\varphi_{n}^{2}(t)}=z\sqrt{\tfrac{1}{2q}}\,\sinh(2\mathcal{J}t)+O\!\left(\tfrac{1}{q^{2}}\right)\,.$ (S5)
Taking the time derivative thus yields
$\frac{dC_{\textrm{\tiny N}}(e^{-iz\mathcal{O}(t)})}{dt}=z\mathcal{J}\sqrt{\tfrac{2}{q}}\,s(t)+\mathcal{O}\!\left(\tfrac{1}{q^{2}}\right)\,.$ (S6)
This result naturally realizes the conjecture proposed in [Susskind:2014jwa, Roberts:2014isa] that the growth rate of quantum complexity is proportional to the operator size.
Appendix B B. Jacobi fields on Berger sphere
In the main text, we analyzed the appearance of conjugate points within the finite-$N$ SYK model under the assumption of homogeneous penalty factors between the odd and even sectors, i.e., eq. (21). This simple choice of penalty factors facilitates the solution of the Jacobi equations on a high-dimensional group manifold. However, it also leads to a trivial evolution of Krylov complexity, since all Krylov basis elements are assigned identical weights.
As solving for the conjugate points with fully arbitrary penalty factors remains analytically challenging, here we consider another simplified model that captures essential features of the problem. Specifically, we focus on a single-qubit system equipped with inhomogeneous penalty factors. Because this system is parameterized by a three-dimensional manifold, the Jacobi equations can be solved exactly, allowing us to determine the location of the conjugate points explicitly. We find that the result agrees with the simplified analysis presented in the main text: the location of the first conjugate point remains bounded from below by $\pi/2$.
For a single qubit, the Nielsen geometry of complexity corresponds to the group manifold $SU(2)$ with a right-invariant metric, often referred to as the Berger sphere [Brown:2019whu]. Despite its simplicity, this setting captures essential features of geodesic deviation and offers exact solutions to the Jacobi equation. We can parametrize $SU(2)$ using Euler angles $\theta_{x,y,z}$ as
$U=e^{i\sigma_{z}\theta_{z}}e^{i\sigma_{y}\theta_{y}}e^{i\sigma_{x}\theta_{x}},$ (S7)
where the Pauli matrices $\sigma_{x,y,z}$ generate the Clifford algebra $\mathbb{CL}_{2}$ via
$\sigma_{x}=\sqrt{2}\psi_{1}\,,\sigma_{y}=\sqrt{2}\psi_{2}\,,\sigma_{z}=-2i\psi_{1}\psi_{2}\,.$ (S8)
This is the simplest algebra realizing the particular algebra (17) by identifying
$X_{\rm o}=\{\sigma_{x},\sigma_{y}\}\,,\qquad X_{\rm e}=\{\sigma_{z}\}\,.$ (S9)
Let us first consider the most general right-invariant metric on this manifold by assigning independent penalty factors $P_{x}$, $P_{y}$, and $P_{z}$ to each direction. Fixing $P_{x}=1$ without loss of generality, the metric takes the form [Brown:2019whu]
$g_{ij}=\left(\begin{array}[]{ccc}\cos^{2}(2\theta_{y})(\cos^{2}(2\theta_{z})+P_{y}\sin^{2}(2\theta_{z}))+P_{z}\sin^{2}(2\theta_{y})&\frac{1}{2}\cos(2\theta_{y})\sin(4\theta_{z})(1-P_{y})&2P_{z}\cos\theta_{y}\sin\theta_{y}\\ \frac{1}{2}\cos(2\theta_{y})\sin(4\theta_{z})(1-P_{y})&\frac{1}{2}(P_{y}+1+(P_{y}-1)\cos(4\theta_{z}))&0\\ 2P_{z}\cos\theta_{y}\sin\theta_{y}&0&P_{z}\end{array}\right).$ (S10)
For the simplest case with $P_{x}=P_{y}=P_{z}$, this gives rise to
$\left.ds^{2}\right|_{P_{x}=P_{y}=P_{z}}=\cos^{2}(2\theta_{y})d\theta_{x}^{2}+d\theta_{y}^{2}+\left(d\theta_{z}+\sin(2\theta_{y})d\theta_{x}\right)^{2}\,.$ (S11)
Using the coordinate transformation
$\phi=\theta_{x}-\theta_{z}\,,\quad\cos\psi=\cos\theta_{y}\cos(\theta_{x}+\theta_{y})\,,\quad\tan\theta=\frac{\sin\theta_{y}}{\cos\theta_{y}\sin(\theta_{x}+\theta_{z})}\,$, (S12)
we can find that the corresponding complexity geometry is nothing but the standard round metric on a three-sphere:
$\left.ds^{2}\right|_{P_{x}=P_{y}=P_{z}}=d\psi^{2}+\sin^{2}\psi\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\,.$ (S13)
Figure 3: Jacobi fields and conjugate points on Berger sphere with odd generators $P_{x}=P_{y}=1$ and different values of the even generator $P_{z}$. Conjugate points appear when $\det\mathcal{J}=0$.
In general, to mimic the straight-line curve generated by a constant control Hamiltonian, we study the simple curve parametrized by
$\theta_{x}(s)=s\,,\quad\theta_{y}(s)=\theta_{z}(s)=0\,$, (S14)
which corresponds to a straight line in the $\sigma_{x}$ direction. The geodesic deviation equation, also known as Jacobi equation,
$\frac{D^{2}V^{\mu}}{ds^{2}}-R^{\mu}{}_{\alpha\beta\sigma}T^{\alpha}T^{\beta}V^{\sigma}=0,$ (S15)
governs the evolution of perturbations $V^{\mu}(s)=(V^{x},V^{y},V^{z})$ along this geodesic. Imposing the initial condition $V^{\mu}(0)=0$, the Jacobi equation can be solved analytically to yield
[…]
where $A=\sqrt{(P_{y}-1)(P_{z}-1)/(P_{y}P_{z})}$. Conjugate points arise when a geodesic ceases to be locally minimizing. Mathematically, they correspond to the appearance of nontrivial Jacobi fields, i.e., solutions to the geodesic deviation equation that vanish at both endpoints. As a result, the existence of a conjugate point corresponds to degeneracies of the Jacobi field, which is determined by the following condition
$\mathcal{J}\begin{bmatrix}C_{1}\\ C_{2}\\ C_{3}\end{bmatrix}=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix},\quad\mathcal{J}=\begin{bmatrix}s&0&0\\ 0&P_{y}\left(\sin(2s)-\sqrt{\frac{P_{z}(P_{y}-1)}{P_{y}(P_{z}-1)}}\sin(2As)\right)&P_{z}(\cos(2As)-\cos(2s))\\ 0&P_{y}(\cos(2s)-\cos(2As))&P_{z}\left(\sin(2s)-\sqrt{\frac{P_{y}(P_{z}-1)}{P_{z}(P_{y}-1)}}\sin(2As)\right)\end{bmatrix}\,$. (S17)
For general $P_{y}$ and $P_{z}$, the determinant of $\mathcal{J}$ is given by
$\det\mathcal{J}(s)=\frac{s}{2}P_{y}P_{z}\left\{(2+B)\cos[2(A-1)s]+(2-B)\cos[2(A+1)s]-4\right\}\,$, (S18)
where $B=(2P_{y}P_{z}-P_{y}-P_{z})/\sqrt{P_{y}P_{z}(P_{y}-1)(P_{z}-1)}$. The conjugate points are thus located at the zeros of the determinant of the Jacobi matrix, i.e.,
$\text{Conjugate point:}\qquad\theta_{x}=\theta_{x}(s_{\rm{conj}})=s_{\rm{conj}}\,,\qquad\text{with}\qquad\det\mathcal{J}(s_{\rm{conj}})=0\,.$ (S19)
Fig. 3 and Fig. 4 show the locations of conjugate points for different penalty factors.
Figure 4: The location of the first conjugate point on Berger spheres (with $P_{x}=1$).
We first reinterpret the numerical results in the main text using the single-qubit model with a homogeneous penalty factor in the odd sector,
$P_{x}=P_{y}=1\,.$ (S20)
Substituting $P_{y}=1$ into eq. (S18) reduces the determinant $\det\mathcal{J}$ to
$\det\mathcal{J}(s)\big|_{P_{x}=P_{y}}\propto\sin s\left[P_{z}\sin s-(P_{z}-1)s\cos s\right]\,.$ (S21)
It is obvious that this expression always vanishes at $s=\pi$, corresponding to conjugate points located at antipodal points (north and south poles) on the sphere. Intuitively, decreasing the penalty factor in the even sector (i.e., $P_{z}<1$) introduces additional conjugate points, as illustrated in Fig. 3. As expected, decreasing the relative ratio $P_{z}$ of even and odd penalty factors moves the conjugate points closer to the initial point. Nevertheless, even in the limiting case $P_{z}\to 0$, the nearest conjugate point remains located at
$\theta_{x}(s_{\ast})\big|_{P_{z}\to 0}=\frac{\pi}{2}\,.$ (S22)
This behavior admits a clear geometric interpretation. In the limit $P_{z}\to 0$, the effective metric reduces to
$\left.ds^{2}\right|_{P_{x}=P_{y}=1,P_{z}\rightarrow 0}=\cos^{2}\left(2\theta_{y}\right)d\theta_{x}^{2}+d\theta_{y}^{2}+\mathcal{O}\left(P_{z}\right)\,.$ (S23)
This is nothing but a two-sphere parameterized by angular coordinates $(2\theta_{x}\,,2\theta_{y})$. As a result, the antipodal points are located at $\theta_{x}=\pi/2$ and $\theta_{x}=\pi$.
We now return to the more general case with inhomogeneous penalties in the odd sector, i.e., $P_{y}\neq P_{x}$. As discussed in the main text, a key difference from the homogeneous setup is that the penalty in the even sector should be taken to infinity for a straight-line trajectory generated by a constant control Hamiltonian to be a geodesic. In the single-qubit model, this limit corresponds to $P_{z}\gg P_{x},P_{y}$. Taking the limit $P_{z}/P_{x,y}\to\infty$ yields a two-frequency transcendental equation
$\cos(2as_{\rm conj})\cos(2s_{\rm conj})+b\sin(2as_{\rm conj})\sin(2s_{\rm conj})=1\,,\quad\text{with}\quad a=\sqrt{\frac{P_{y}-1}{P_{y}}}\,,b=\frac{2P_{y}-1}{2\sqrt{P_{y}(P_{y}-1)}}\,$, (S24)
which generally lacks a simple, closed-form algebraic solution. Instead, Fig. 4 shows the numerical solution for the location of the first conjugate point. We find that the nearest conjugate point remains at $\theta_{x}(s_{\ast})\geq\pi/2$, with saturation only in the limit $P_{y}\to 0$. To gain analytic control over this solution, we focus on the regime $P_{y}\ll P_{x}=1$. Expanding the analytic expression for $\det\mathcal{J}$ in this regime yields
$\lim_{P_{z}\to\infty}\det\mathcal{J}(s)\propto-2\sqrt{P_{y}}\sin(2s)\sinh\left(\frac{2s}{\sqrt{P_{y}}}\right)+4P_{y}\cos(2s)\cosh\left(\frac{2s}{\sqrt{P_{y}}}\right)+\mathcal{O}(P_{y})\,.$ (S25)
A straightforward asymptotic analysis then shows that the first conjugate point occurs at
$s_{\ast}\approx\frac{\pi}{2}+\sqrt{P_{y}}\,$, (S26)
showing that the nearest conjugate point is located at $s_{\ast}\approx\frac{\pi}{2}$ in the limit $P_{y}\to 0$.
In summary, the single-qubit model demonstrates that the first conjugate point associated with the straight-line geodesic generated by a constant control Hamiltonian (both in the homogeneous case and in the inhomogeneous setup with $G_{\rm even}\gg 1$) is always bounded from below by
$\theta_{x}(s_{\ast})\geq\frac{\pi}{2}\,.$ (S27)
We have shown that this bound also holds for the SYK model under the homogeneous configuration discussed in the main text. Although a rigorous proof for general finite-dimensional systems is still lacking, it is natural to expect that this lower bound persists in more complex models such as SYK, where the operator dynamics exhibit similar geometric features.